OpenStudy (anonymous):

I'm a bit confused about dot products - could someone explain them to me or point me to a place where I could find help? Thanks!

OpenStudy (anonymous):

A dot product is a way of saying multiply the x value(i) and the y values (j) then adding them together. But there are other things you can do with them as well. But the basic Dot Product: v=(V1,V2) W=(W1,W2) so: [(V1)(W1)] + [(V2)(W2)] EX: If v=(1,2) and w=(3,-4) then: [1(3)] + [2(-4)] = 3+(-8) = -5 -5 is the dot product But there are other things to do as well like using the dot product to find projections, cosine angle etc... But that was a basic Dot Product.

OpenStudy (anonymous):

Thank you so much!